OpenStudy (anonymous):

PLEASE HELP!!! 2.5 qt/min = ____ gal/h

OpenStudy (espex):

Set it up so that you can use a dimensional analysis on it. Using your conversions you set it so the units cancel to give you what you're after, then math is all that's left.\[\frac{2.5~qt}{1~min} \dot{}\frac{60~min}{1~hr}\dot{} \frac{1~gal}{4~qt}=\]